OpenStudy (vortish):

a company offers a starting salary of 31,000 a year with raises of 2,000 per year find the total salary over a ten year period

OpenStudy (anonymous):

So you start with 31000 and every year 2000 is added to that. So the salary each year is 31000+2000y where y is the numbr of years if you call the first starting year y=0.

OpenStudy (vortish):

ok i understand that but im still confused by it

OpenStudy (anonymous):

Think of it as a function. Salary per year S(y)=31000+2000y S(0)=31000 S(1)=31000 + 2000*1 S(2)=31000 + 2000*2 Now you add all the values S(0)+S(1)+S(2)...

OpenStudy (vortish):

i come up with 96,000 but that does not seem right

OpenStudy (anonymous):

How did you get that?

OpenStudy (vortish):

doing the way you had it on the board by s or year ten it is 31,000+20000*10 right so in year ten its 51,000 just for that year and if you add up all the other years its more than that right

OpenStudy (anonymous):

Over a ten year period probably means go to y=9 since you started at y=0. So add the amount earned in year 0 to the amount earned in year 1 to...

OpenStudy (vortish):